# A radial scalar product for Kerr quasinormal modes

### v3

@@ Abstract
A scalar product for quasinormal mode solutions to Teukolsky’s homogeneous radial equation is presented. Evaluation of this scalar product can be performed either by direct integration, or by evaluation of a confluent hypergeometric functions. For direct integration, it is explicitly shown that the quasinormal modes’ radial functions are regular on a family of physically bounded complex paths. The related scalar product will be useful for better understanding analytic solutions to Teukolsky’s radial equation, particularly the quasi-normal modes, their potential spatial completeness, and whether the quasi-normal mode overtone excitations may be estimated by spectral decomposition rather than fitting. With that motivation, the scalar product is applied to confluent Heun polynomials where it is used to derive their peculiar orthogonality and eigenvalue properties. A potentially new relationship is derived between the confluent Heun polynomials’ scalar products and eigenvalues. Using these results, it is shown for the first time that Teukolsky’s radial equation (and perhaps similar confluent Heun equations) are, in principle, exactly tri-diagonalizable. To this end, “canonical” confluent Heun polynomials are conjectured.
@@ I Introduction
The purpose of this article is to introduce a framework (via a radial scalar product) that may enable a deeper practical understanding of ringdown’s “overtones”, and perhaps gravitational waves from compact binaries broadly.
@@
The term “ringdown” specifically refers to linear combinations of BH Quasi-Normal Modes (QNMs) Press (1971); Leaver (1985); Press and Teukolsky (1973). Thus the mathematical structure of ringdown is that of the QNMs. Due to their astrophysical relevance, the QNMs of Kerr BHs are perhaps the most studied Berti et al. (2003); Berti and Klein (2014); Berti et al. (2006b); London (2020); London et al. (2014); London (2023); London and Fauchon-Jones (2019); Husa et al. (2016); Hamilton et al. (2021, 2023); Cotesta et al. (2020); Pompili et al. (2023); Ossokine et al. (2020); Hughes et al. (2019); Ghosh et al. (2018); Yang et al. (2012, 2013); Zimmerman and Mark (2016); Lagos and Hui (2022).
@@
However, in 1985 Leaver presented what was perhaps the first Frobenius-type series solution for the QNM’s radial functions Leaver (1985). Leaver later pointed out that, although such series solutions may be locally convergent, they may diverge at either the event horizon or at spatial infinity Leaver (1986); Ronveaux and Arscott (1995). It wasn’t until 1996 that Mano, Suzuki and Takasugi (MST) presented a kind of “trick” that allows for the analytic matching between one series expansion valid at the BH horizon, and another valid near spatial infinity Mano et al. (1996); Sasaki and Tagoshi (2003). Mostly used in the context of gravitational wave self-force and particle perturbation theory, various developments upon the MST method have been presented Isoyama et al. (2022); Casals and Zimmerman (2019); Sasaki and Tagoshi (2003); Fujita and Tagoshi (2004).
In parallel, there has been broad and growing interest in a generalization of the hypergeometric equation called the Heun equation Ronveaux and Arscott (1995); Chen and Jing (2023); Fiziev (2009a); Hortacsu (2012); Dariescu et al. (2021); Magnus et al. (2021); Bonelli:2021uvf. Unlike the hypergeometric equation, which has three regular singular points, the Heun equation has four Ronveaux and Arscott (1995). It was perhaps first pointed out by Leaver and then Fiziev that Teukolsky’s radial equation is an instance of the confluent Heun equation, also known as a generalized spheroidal equation Fiziev (2009a). Fiziev’s work appears to coincide with what is perhaps a recent and multi-disciplinary interest in Heun equationsHortacsu (2012). For example, it has very recently been observed that a careful consideration of the confluent Heun function’s asymptotic behavior allow them to be applied directly to astrophysically relevant scenarios in particle perturbation theory (e.g. gravitational waves from extreme mass ratio inspirals) Chen and Jing (2023).
@@
    The properties of many special functions in physics may be framed using ideas from Sturm-Liouville theory and functional analysis Arfken et al. (2013); Kristensson (2010); Courant and Hilbert. (1954); Adkins and Weintraub (2012). There, solutions to differential equations are conceptualized as eigenvectors, where the vector scalar products (i.e. symmetric nondegenerate bilinear forms Adkins and Weintraub (2012)) are defined by an integral whose kernel is weighted such that the differential equations is formally self-adjoint. For example, the scalar product for the QNM’s angular functions can be framed as that of the Jacobi polynomials (i.e. with the Jacobi weight function), and the Jacobi differential equation is self-adjoint with respect to that scalar product Fackerell and Crossman (2008). In analogy:
@@
Regarding (Q1), many studies have encountered the need to perform integrals over solutions to Teukolsky’s radial equation, e.g. Green et al. (2023); Motl and Neitzke (2003); Hughes (2000a); Leaver (1986); Andersson (1997); Jaramillo et al. (2021). Perhaps most relevant to the present article, Leaver (Ref. Leaver (1986)) studied Schwarzschild (i.e. non-spinning) BHs, and used a Green’s function approach to derive various spectral-type expansions, including those for QNMs. Similar approaches have recently been used to study non-linear (particularly quadratic) QNMs of Schwarzschild, e.g. Lagos and Hui (2022). This branch of research intersects with (Q1), however it does not directly draw from Sturm-Liouville theory due to apparent divergences in QNM-type problems that are not present in classical resonance problems [^3] Similarly, Ref. Nollert and Price (1999) implies that the QNM problem may not be compatible with Sturm-Liouville theory because QNMs damp in the time domain. In Sec. IV of this article, this is shown to not be the case. Leaver (1986). Interestingly, Ref. Leaver (1986) observes that exactly such divergence may be overcome by allowing for a radial integration path that is complex valued. In this article, the same idea is encountered for Kerr, and is motivated by coordinate dependence.
@@
In an adjacent branch of research, many studies have made use of what’s called the “formal adjoint” of operators related to QNMs where, in effect, a weight function of $1$ is considered, e.g. Teukolsky and Press (1974); Wald (1979); Green et al. (2023); Dias et al. (2022); Fiziev (2009a); Jaramillo et al. (2021). From that perspective, Teukolsky’s radial equation is not self-adjoint due to the weight function or scalar product used. Consequently, it appears that an investigation of Teukolsky’s radial equation in the context of Sturm-Liouville theory with a non-trivial weight has not yet been pursued.
@@
Regarding (Q2), a number studies have attempted to use classical polynomials to represent solutions to the QNM’s radial functions, e.g. Chung et al. (2023a, b); Ghojogh et al. (2023). These efforts encounter the following mathematical difficulty: the asymptotic (i.e. large distance from the BH) behavior of the classical polynomials is incompatible with the QNM’s radial functions, specifically the constraint that gravitational radiation from QNMs must be purely outgoing towards spatial infinity Teukolsky (1973); Press and Teukolsky (1973); Teukolsky and Press (1974). As a result, one encounters a “generalized spectral problem”, which may result in the appearance of non-physical solutions Chung et al. (2023a). In contrast, a purely spectral formulation would result in a matrix representation of the problem that is e.g. tridiagonal, or generally band-diagonal, and has no artificial eigensolutions Cook and Zalutskiy (2014); Fackerell and Crossman (2008). Although it has been recognized that Teukolsky’s radial equation is a confluent Heun equation, it appears that confluent Heun polynomials have not been used to frame the QNM problem as purely spectral (until this work).
@@
In 2014, Cook and Zalutskiy studied the connection between “algebraically special” QNMs of nonspinning BH’s, and “total transmission” QNMs in the presence of BH spin Cook and Zalutskiy (2014). There, it was found that for special QNM frequencies, the associated radial functions can be directly identified with select confluent Heun polynomials. With (Q2) in mind, there is good reason to ask whether the confluent Heun polynomials are only relevant to the total transmission QNMs, or whether there exists a single formulation that connects confluent Heun polynomials to all QNM solutions?
@@ II.1 Scope
This article discusses QNM solutions to Teukolsky radial equation in the context of isolated Kerr BHs, absent any effects beyond GR. Some aspects of the discussion make minimal physical assumptions, while others are specific to gravitational wave ringdown. Broadly, the reader is cautioned that the author does not aspire to the language of formal mathematical rigor (e.g. the text will not include theorems and lemmas).
Most mathematical arguments make no assumption about spin weight, $s$. However, $s\in\{-2,+2\}$ are given particular attention due to their correspondence with gravitational radiation. Further, Teukolsky’s radial equation, even when formatted for the QNMs, may be considered for any frequency. Therefore, much of this article’s content is not specific to the QNM’s frequencies, but rather the structure of Teukolsky’s radial equation when formatted with QNM boundary conditions.
This article’s central result is a frequency dependent scalar product; however, it is also noted in Sec. IV.2 that this is but one of a number of equivalent formulations. For example, a frequency independent scalar product is shown in Eq. (31), and its relation to this article’s frequency dependent scalar product is described in adjacent text. That scalar product may be of future interest to an investigation analogous to Ref. London (2023), where different QNM frequencies are studied simultaneously to learn about the completeness and biorthogonality properties of the QNM’s spheroidal harmonics.
@@ III Teukolsky’s equation with QNM boundary conditions
In this section we connect the observable gravitational wave polarizations with this article’s topical starting point, Teukolsky’s radial equation with QNM boundary conditions. Along the way, we will encounter spin weighted fields Newman and Penrose (1966, 1962), the definition of QNMs Nollert (1999); Berti et al. (2016), the importance of a scalar product and polynomials to Teukolsky’s angular equation, and how QNM solutions to Teukolsky’s radial equation may be framed in close analogy with the angular case Leaver (1985). These ideas converge on Teukolsky’s radial equation in a simplified form Cook (2020); Fiziev (2010); Chen and Jing (2023). That equation sets the stage for our application of Sturm-Liouville theory Pinchover and Rubinstein (2005); Abramowitz and Stegun (1964); Arfken et al. (2013); Courant and Hilbert. (1954), and the development of a radial scalar product in Sec. IV.
@@ III.1 From gravitational wave strain to the perturbative wave quantity
[…]
In Eq. (1a), $h$ is the complex valued gravitational wave strain Ruiz et al. (2008). We will consider it to depend on Boyer–Lindquist coordinates, where time, radius, polar angle and azimuthal angle are denotes by $t$, $r$, $\theta$ and $\phi$ respectively Boyer and Lindquist (2004). In Eq. (1b), $\psi_{4}$ is the $5$’th Weyl scalar in Newman-Penrose notation Ruiz et al. (2008); Newman and Penrose (1962). In that formalism, $\psi_{4}$ is a complex valued scalar field of spin weight $s=-2$ Newman and Penrose (1966); Teukolsky (1973). Henceforth, complex valued qualities encode or modify the relative importance of $h_{+}$ and $h_{\times}$ within $h$. Within the study of gravitationally perturbed isolated BHs, two Weyl scalars are of particular interest: $\psi_{4}$ ($s=-2$), and the closely related $\psi_{0}$ ($s=+2$). Both encode information about gravitational waves Teukolsky (1973).
@@ III.2 Teukolsky’s equation and its QNM solutions
[…]
The fact that $\mathcal{L}_{tr\theta\phi}$ (see Eq. 3) does not explicitly depend on $t$ and $\phi$ means that $\psi$ of the form
@@
may separate [^4] The ansatz provided in Eq. (7) obscures the fact that our net solution will depend on the tortoise coordinate via the retarded time Press and Teukolsky (1973); Teukolsky (1973). The radial part of the retarded time is ultimately encoded within $R(r)$. $\mathcal{L}_{tr\theta\phi}$ into two coupled equations for $R(r)$ and $S(\theta)$. In Eq. (7), $\tilde{\omega}$ is generally a complex valued frequency:
@@
In Eq. (8), one should think of any physical situation in which an isolated BH experiences a momentary and small perturbation. The post-merger of two sufficiently compact objects is an example of particular importance to astrophysics Helfer et al. (2017); Abbott et al. (2021b, a); Baker et al. (2006). For the simplest scenarios, the resulting gravitational radiation will, qualitatively, ring down Nollert (1999); Press (1971); Berti et al. (2007); Kokkotas and Schmidt (1999); Nollert and Price (1999). The central frequency of that ringing will be $\omega$, and the exponential decay rate of that ringing will be $1/\tau$, where $\tau$ is strictly positive; otherwise, Eq. (7) would diverge at late times, and the BH would be linearly unstable Vishveshwara (1970); Press and Teukolsky (1973); Press (1971); Whiting (1989).
In general, gravitational radiation will be a sum of such mode-like terms (Eq. 7) weighted by constant factors that depend on spacetime geometry and the perturber Berti and Cardoso (2006); Kamaretsos et al. (2012); London et al. (2014); London (2020). The net spatial dependence of each term in the sum defines a single QNM Berti et al. (2007); Kokkotas and Schmidt (1999); Nollert and Price (1999). There will always be a period in time for which QNM radiation dominates over other solutions to the homogeneous radial equation, such as power-law tails, prompt emission, and nonlinear QNMs Leaver (1986); Lagos and Hui (2022); Andersson (1997). This inherently poorly defined region of QNM dominance is colloquially referred to as ringdown Press (1971); Nollert (1999).
@@
In Eq. (III.2), $u=\cos(\theta)$, $\mathcal{L}{{}_{u}}$ will be referred to as the angular differential operator, and $\mathcal{L}{{}_{r}}$ will be referred to as the radial differential operator. The angular functions, $S(u)$, are called spheroidal harmonics, and we will refer to $R(r)$ as Teukolsky’s radial functions. Both are known to be closely related to the confluent Heun functions [^5] As noted by Leaver, they are also called generalized spheroidal wave functions Leaver (1986). Fiziev (2010); Cook and Zalutskiy (2014); Cho et al. (2009); Chen and Jing (2023).

### v4

@@ Abstract
A scalar product for quasinormal mode solutions to Teukolsky’s homogeneous radial equation is presented. Evaluation of this scalar product can be performed either by direct integration, or by evaluation of a confluent hypergeometric functions. The related scalar product will be useful for better understanding analytic solutions to Teukolsky’s radial equation, particularly the quasi-normal modes, their potential spatial completeness, and whether the quasi-normal mode overtone excitations may be estimated by spectral decomposition rather than fitting. With that motivation, the scalar product is applied to confluent Heun polynomials where it is used to derive their peculiar orthogonality and eigenvalue properties. A potentially new relationship is derived between the confluent Heun polynomials’ scalar products and eigenvalues. Using these results, it is shown for the first time that Teukolsky’s radial equation (and perhaps similar confluent Heun equations) are, in principle, exactly tri-diagonalizable. To this end, “canonical” confluent Heun polynomials are conjectured.
@@ I Introduction
The purpose of this article is to introduce a framework via a radial scalar-product that may offer a deeper, practical understanding of ringdown “overtones,” and potentially of gravitational waves from compact binaries more broadly. The primary technical focus of this work is the vector space structure of the eigensolutions to Teukolskys equation for Quasi-Normal Modes (QNMs), considered after the imposition of physical boundary conditions.
@@
Here, ringdown will refer to linear combinations of BH QNMs Press (1971); Leaver (1985); Press and Teukolsky (1973). While the gravitational waves from BBH postmerger will refer to all linear and nonlinear contributions London et al. (2014), the mathematical structure of ringdown is that of the QNMs. Due to their astrophysical relevance, the QNMs of Kerr BHs are perhaps the most studied Berti et al. (2003); Berti and Klein (2014); Berti et al. (2006b); London (2020); London et al. (2014); London (2023); London and Fauchon-Jones (2019); Husa et al. (2016); Hamilton et al. (2021, 2023); Cotesta et al. (2020); Pompili et al. (2023); Ossokine et al. (2020); Hughes et al. (2019); Ghosh et al. (2018); Yang et al. (2012, 2013); Zimmerman and Mark (2016); Lagos and Hui (2022).
@@
However, in 1985 Leaver presented what was perhaps the first Frobenius-type series solution for the QNM’s radial functions Leaver (1985, 1986). Leaver’s work on QNMs intersects with topics in gravitational wave self-force, in which asymptotic solutions to Teukolskys radial equation play a key role Isoyama et al. (2022); Casals and Zimmerman (2019); Sasaki and Tagoshi (2003); Fujita and Tagoshi (2004).
In parallel, there has been broad and growing interest in a generalization of the hypergeometric equation called the Heun equation Ronveaux and Arscott (1995); Chen and Jing (2023); Fiziev (2009a); Zarrinkamar and Hortaçsu (2018); Dariescu et al. (2021); Magnus et al. (2021); Bonelli et al. (2022). Unlike the hypergeometric equation, which has three regular singular points, the Heun equation has four Ronveaux and Arscott (1995). It was perhaps first pointed out by Leaver and then Fiziev that Teukolsky’s radial equation is an instance of the confluent Heun equation, also known as a generalized spheroidal equation Fiziev (2009a). Fiziev’s work appears to coincide with what is perhaps a recent and multi-disciplinary interest in Heun equationsZarrinkamar and Hortaçsu (2018). For example, it has very recently been observed that a careful consideration of the confluent Heun function’s asymptotic behavior allow them to be applied directly to astrophysically relevant scenarios in particle perturbation theory (e.g. gravitational waves from extreme mass ratio inspirals) Chen and Jing (2023).
@@
    The properties of many special functions in physics may be framed using ideas from Sturm-Liouville theory and functional analysis Arfken et al. (2013); Kristensson (2010); Courant and Hilbert (1989); Adkins and Weintraub (2012). There, solutions to differential equations are conceptualized as eigenvectors, where the vector scalar products (i.e. symmetric nondegenerate bilinear forms Adkins and Weintraub (2012)) are defined by an integral whose integrand is weighted such that the differential equations is formally self-adjoint. For example, the scalar product for the QNM’s angular functions can be framed as that of the Jacobi polynomials (i.e. with the Jacobi weight function), and the Jacobi differential equation is self-adjoint with respect to that scalar product Fackerell and Crossman (2008). Further, the Jacobi polynomials are compatible with the spheroidal harmonic’s analytic (i.e. regular and differentiable) structure Fackerell and Crossman (2008). In analogy:
@@
Regarding (Q1), many studies have encountered the need to perform integrals over solutions to Teukolsky’s radial equation, e.g. Green et al. (2023); Motl and Neitzke (2003); Hughes (2000a); Leaver (1986); Andersson (1997); Jaramillo et al. (2021). Perhaps most relevant to the present article, Leaver (Ref. Leaver (1986)) studied Schwarzschild (i.e. non-spinning) BHs, and used a Green’s function approach to derive various spectral-type expansions, including those for QNMs. Similar approaches have recently been used to study non-linear (particularly quadratic) QNMs of Schwarzschild, e.g. Lagos and Hui (2022). This branch of research intersects with (Q1), however it does not directly draw from Sturm-Liouville theory due to apparent divergences in QNM-type problems that are not present in classical resonance problems [^3] Similarly, Ref. Nollert and Price (1999a) implies that the QNM problem may not be compatible with Sturm-Liouville theory because QNMs damp in the time domain. In Sec. IV of this article, this is shown to not be the case. In paricular, the QNM problem is non-Hermitian; however, this does not preclude the self-adjoint perspective taken in this work. Leaver (1986). Interestingly, Ref. Leaver (1986) observes that exactly such divergence may be overcome by allowing for a radial integration path that is complex valued. In this article, the same idea is encountered for Kerr, and is motivated by coordinate dependence.
@@
In an adjacent branch of research, many studies have made use of what’s called the “formal adjoint” of operators related to QNMs where, in effect, a weight function of $1$ is considered, e.g. Teukolsky and Press (1974); Wald (1979); Green et al. (2023); Dias et al. (2022); Fiziev (2009a); Jaramillo et al. (2021). From that perspective, Teukolsky’s radial equation is not self-adjoint due to the weight function or scalar product used. Consequently, it appears that an investigation of the QNMs’ radial functions in the context of Sturm-Liouville theory with a physically motivated weight has not yet been pursued.
@@
Regarding (Q2), a number studies have attempted to use classical polynomials to represent solutions to the QNMs’ radial functions, e.g. Chung et al. (2023a, b); Ghojogh et al. (2023). These efforts encounter the following mathematical difficulty: while Teukolsky’s radial equation may be similarity transformed to be similar to the equation that governs e.g. Jacobi polynomials, the related transformations differ from those required by the physical boundary conditions Teukolsky (1973); Press and Teukolsky (1973); Teukolsky and Press (1974). In turn, this means that unlike Teukolsky’s angular eigenvalue problem for QNMs, Teukolsky’s radial eigenvalue problem cannot be exactly represented in terms of a standard spectral problem involving classical polynomials. Instead, one encounters a “generalized spectral problem” Chung et al. (2023a); Ghojogh et al. (2023). Although it has been recognized that Teukolsky’s radial equation is a confluent Heun equation, it appears that confluent Heun polynomials have not been used to frame the QNM radial problem as purely spectral until the present work.
@@
In 2014, Cook and Zalutskiy studied the connection between confluent Heun polynomails, “algebraically special” QNMs of nonspinning BH’s, and “total transmission” QNMs in the presence of BH spin Cook and Zalutskiy (2014). There, it was found that for special QNM frequencies, the associated radial functions can be directly identified with select confluent Heun polynomials. With (Q2) in mind, there is good reason to ask whether the confluent Heun polynomials are only relevant to the total transmission QNMs, or whether there exists a single formulation that connects confluent Heun polynomials to all QNM solutions?
@@ II.1 Scope
This article discusses QNM solutions to Teukolsky radial equation in the context of isolated Kerr BHs, absent any effects beyond GR. Some aspects of the discussion make minimal physical assumptions, while others are specific to gravitational wave ringdown. Broadly, this work studies the discrete spectra of Teukolsky’s radial equation. By construction, all versions of this equation are predicated on a single QNM frequency (i.e. the radial equation for QNMs may only be defined for a single frequency at a time).
The underlying perspective of this work is that the structure QNM frequencies is influenced by the underlying structure of the discrete spectra of the radial equation for single frequencies. This concept is outlined in Sec. III.1, and it is encapsulated by the overtone number, $n$. It is known that many references refer to the overtone number as being heuristically related to an “apparent” order in the imaginary part of QNM frequencies Berti et al. (2006a); Nollert (1999); Nollert and Price (1999b). That apparent order is equally present in the QNM separation constants, and in the context of this work, each QNM separation constant is one of many radial eigenvalues.
Most mathematical arguments make no assumption about spin weight, $s$. However, $s\in\{-2,+2\}$ are given particular attention due to their correspondence with gravitational radiation. Further, Teukolsky’s radial equation, even when formatted for the QNMs, may be considered for any frequency. Therefore, much of this article’s content is not specific to the QNM’s frequencies, but rather the structure of Teukolsky’s radial equation when formatted with QNM boundary conditions. The reader is cautioned that the author does not aspire to the language of formal mathematical rigor (e.g. the text will not include theorems and lemmas).
This article’s primary focus is a frequency dependent scalar product that is uniquely determined by a choice of radial coordinate and the QNM boundary conditions. While it is noted in Sec. IV.2 that this product is related to a range of ostensibly frequency independent products, a key feature of all such products is that they are conditioned on the frequency parameter used to separate Teukolsky’s equation. In that sense, all such products are frequency dependent: some place the frequency dependence in a pre-factor in front of the QNM radial functions, while others, including the one studies here, place the frequency dependence within the scalar product’s weight function. This concept is expanded upon in Sec. IV.2.
@@ III Teukolsky’s equation with QNM boundary conditions
In this section we connect the observable gravitational wave polarizations with this article’s topical starting point, Teukolsky’s radial equation with QNM boundary conditions. Along the way, we will encounter spin weighted fields Newman and Penrose (1966, 1962), the definition of QNMs Nollert (1999); Berti et al. (2016), the importance of a scalar product and polynomials to Teukolsky’s angular equation, and how QNM solutions to Teukolsky’s radial equation may be framed in close analogy with the angular case Leaver (1985). These ideas converge on Teukolsky’s radial equation in a simplified form Cook (2020); Fiziev (2010); Chen and Jing (2023). That equation sets the stage for our application of Sturm-Liouville theory Pinchover and Rubinstein (2005); Abramowitz and Stegun (1964); Arfken et al. (2013); Courant and Hilbert (1989), and the development of a radial scalar product in Sec. IV.
@@ III.1 From gravitational wave strain to the perturbative wave quantity
$=\;\int_{-\infty}^{t}\int_{-\infty}^{t^{\prime}}\psi_{4}(t^{\prime\prime},r,\theta,\phi)\;dt^{\prime\prime}dt^{\prime}\;.$ (1b)
In Eq. (1a), $h$ is the complex valued gravitational wave strain Ruiz et al. (2008). We will consider it to depend on Boyer–Lindquist coordinates, where time, radius, polar angle and azimuthal angle are denotes by $t$, $r$, $\theta$ and $\phi$ respectively Boyer and Lindquist (2004). In Eq. (1b), $\psi_{4}$ is the $5$’th Weyl scalar in Newman-Penrose notation Ruiz et al. (2008); Newman and Penrose (1962). In that formalism, $\psi_{4}$ is a complex valued scalar field of spin weight s=2s=-2 Newman and Penrose (1966); Teukolsky (1973). Henceforth, complex valued qualities encode or modify the relative importance of $h_{+}$ and $h_{\times}$ within $h$. Within the study of gravitationally perturbed isolated BHs, two Weyl scalars are of particular interest: $\psi_{4}$ (s=2s=-2), and the closely related $\psi_{0}$ ($s=+2$). Both encode information about gravitational waves Teukolsky (1973).
@@ III.2 Teukolsky’s equation and its QNM solutions
$T(t,r,\theta,\phi)\;=\;0\;.$ (6)
The structure of $\mathcal{L}_{tr\theta\phi}$ (see Eq. 3) is known to be such that $\psi$ of the form
@@
separates [^4] The ansatz provided in Eq. (7) obscures the fact that our net solution will depend on the tortoise coordinate via the retarded time Press and Teukolsky (1973); Teukolsky (1973). The radial part of the retarded time is ultimately encoded within $R(r)$. $\mathcal{L}_{tr\theta\phi}$ into two coupled equations: one for $R(r)$ and another for $S(\theta)$. In Eq. (7), $\tilde{\omega}$ is generally a complex valued frequency:
@@
In Eq. (8), one should think of any physical situation in which an isolated BH experiences a momentary and small perturbation. The post-merger of two sufficiently compact objects is an example of particular importance to astrophysics Helfer et al. (2017); Abbott et al. (2021b, a); Baker et al. (2006). For the simplest scenarios, the resulting gravitational radiation will, qualitatively, ring down Nollert (1999); Press (1971); Berti et al. (2007); Kokkotas and Schmidt (1999); Nollert and Price (1999a). The central frequency of that ringing will be $\omega$, and the exponential decay rate of that ringing will be $1/\tau$, where $\tau$ is strictly positive; otherwise, Eq. (7) would diverge at late times, and the BH would be linearly unstable Vishveshwara (1970); Press and Teukolsky (1973); Press (1971); Whiting (1989).
In general, gravitational radiation will be a sum of such mode-like terms (Eq. 7) weighted by constant factors that depend on spacetime geometry and the perturber Berti and Cardoso (2006); Kamaretsos et al. (2012); London et al. (2014); London (2020). The net spatial dependence of each term in the sum defines a single QNM Berti et al. (2007); Kokkotas and Schmidt (1999); Nollert and Price (1999a). There will always be a period in time for which QNM radiation dominates over other solutions to the homogeneous radial equation, such as power-law tails, prompt emission, and nonlinear QNMs Leaver (1986); Lagos and Hui (2022); Andersson (1997). This inherently poorly defined region of QNM dominance is colloquially referred to as ringdown Press (1971); Nollert (1999).
@@
In Eq. (9), $\cos(\theta), u\mathcal{L}{{}_{u}}$ will be referred to as the angular differential operator, and $\mathcal{L}{{}_{r}}$ will be referred to as the radial differential operator. The angular functions, $S(u)$, are called spheroidal harmonics, and we will refer to $R(r)$ as Teukolsky’s radial functions. Both are known to be closely related to the confluent Heun functions [^5] As noted by Leaver, they are also called generalized spheroidal wave functions Leaver (1986). Fiziev (2010); Cook and Zalutskiy (2014); Cho et al. (2009); Chen and Jing (2023).